OpenStudy (anonymous):

Approximately 9% of high school athletes go on to play sports in college. Of these college athletes, only 1.3% go on to play professional sports. What is the probability that a high school athlete will go on be a professional athlete? a)0.04% b)0.14% c)0.117% d)0.22%

OpenStudy (ddcamp):

P(A and B) = P(A) * P(B) P(College then Proffesional) = P(college) * P(proffesional) = 0.09 * 0.013 = 0.00117 or 0.117%